# Optimizing Continuous-Wave Pumped Entanglement-based QKD in Noisy Environment

### v1

@@ II.2 Coincidence Measurement in Noisy Settings
    The height of the cross-correlation function is reduced as the noise increases
@@
where $B$ is the number of pairs generated, $\eta$ is the detection efficiency. $\eta$ is the combination of all loss parameters. With this, the probability of simultaneously arriving photons in both detectors can be estimated with
@@
$S_{B}=B\eta_{A}\eta_{B}+N_{B}+DC_{B}$ (4)
where $N_{A}$ and $N_{B}$ are the count of noise entered in detector $A$ and detector $B$, $DC_{A}$ and $DC_{B}$ are the dark count rates of detector $A$ and detector $B$. In the measurement, there is a nonzero probability that two detection events fall within the same time window, even though they are not from the SPDC process. This is known as accidental coincidence events. These events are the primary source of noise in coincidence-based measurement systems. The chance of accidental coincidence increases when noise photons probabilistically fall in the time-window measurement together. Therefore, the accidental coincidence is calculated by multiplying single events from two detectors with the chosen coincidence time window which can be written as
@@
where $\sigma$ is the detector jitter and $\tau$ is the measurement time. This method can be applied in numerical simulations to investigate quantum state tomography of entangled qubits with measurements affected by time uncertainty [27]. However, this approximation fails when systems operate in extreme noise regimes. Also, this approximation does not allow one to study individual detector characteristics. In our experiment, where only one detector receives noise-induced clicks, the probabilistic distribution of the coincidence events does not show a clean Gaussian profile, rendering a normal distribution that does not fit our collected data. Neumann et al. [20] investigated coincidence events using the normal distribution function for various coincidence values, accounting for variations in the Full Width at Half Maximum (FWHM) and shifts caused by dispersion in optical fibers. Using the detector function of the form of Eq. (6), they characterized the changes in the detector parameters using linear functions. In contrast, our observations reveal a polynomial growth in FWHM and a shift in the coincidence time. This discrepancy arises because previous work assumed equal photon pair detection rates across all detectors in a low-noise scenario. Consequently, FWHM changes and coincidence peak shifts appeared linear. Our results show that noise can shift detection timing, particularly when detection rates are imbalanced between detectors. To address this, we should be able to cater to the changes in individual detectors separately. By incorporating noise-induced variation parameter with exponential distribution coincidence histogram, we could detect changes in the individual detectors as follows
@@
where $W_{s(i)}$ is a coefficient used to accommodate the change in the width of coincidence events, $t_{d}$ is the dead time of the detector and $C$ is the centroid value of the coincidence histogram. $1/W$ can be used to measure the FWHM in the time domain. By convoluting detector function in Eq. (6) with $A_{i}$ and $A_{j}$ we can obtain a complete coincidence histogram function of the form:
@@
where $\sigma_{s}$ and $\sigma_{i}$ are the timing uncertainty (jitter) of the detectors used at signal and idler arms, respectively, $t_{D}$ is the shift of centroid point of coincidence histogram from the origin, and “$\mathrm{erfc}$” represents the complimentary error function. Detector timing jitter is predominantly responsible for the width of the coincidence events. The timing uncertainty of the detectors (jitter) is responsible for the width of the coincidence histogram plots. If the temporal coherence of the collected photons is significantly greater than the jitter of the detection system, the measured spectral profile will reflect the true spectral characteristics of the photons more closely [28, 29]. The coherence time of our photon pairs is about 21.5 fs (corresponding to the spectral bandwidth of 30 nm), which is negligible and overshadowed by the timing jitter of the detectors. Therefore, timing uncertainty due to this factor could be omitted. Additionally, the temporal resolution of the timestamp unit also contributes to the width of coincidence events if it is greater than the timing jitter. Since the temporal resolution of our timestamp unit is 81 ps, we can omit this factor as well. When the noise is low, we can approximate $W_{s(i)}=\sigma_{s(i)}$. As noise increases, we adjusted the value of $W_{s(i)}$ to fit with the experimental data. Theoretical fit for the collected experimental data is shown in Fig.  2(a). The change in the FWHM in different noise settings is shown in Fig. 2(b). We observe that $W_{s(i)}$ is proportional to the FWHM of the coincidence histogram. The experimentally obtained shift of the start-stop histogram was directly incorporated into Eq. (9) during data fitting. One parameter that largely depends on the shape of the coincidence histogram is the variation in the dead time of the detectors due to high photon detection events. We have investigated the change in dead time with increased event rates in the detector. We used experimentally obtained $t_{d}$ values (discussed in Section II.3) in our model.
@@ II.3 Characterization of Detector Parameters
The SPCM-AQRH-10 detector is a commonly used single-photon detection module in QKD. This is an actively quenched Geiger mode detector, and the manufacturer-provided parameters of this detector are summarized in Table. 1. To study the count rate-dependent detector parameters, from the experimental setup described in Fig. 1, we only considered the APD2 output and investigated detector parameters such as detector dead time and twilight events. For this, the singles values recorded from APD2 were subtracted from their neighboring photon time values, that is, $\Delta t=t_{n+1}-t_{n}$ where $t$ is the time of arrival of photons and $n$ is the index value. The number of measured time separations ($\Delta$) is plotted as a histogram in Fig. 3. This figure gives interesting results about many detector parameters, including dead time and twilight events. The dead time in single-photon avalanche photo detectors [21] is often described as a temporal window in which photons cannot be detected, which means that the dead time can be estimated by knowing the separation from 0 to the offset value of $\Delta t$ time in Fig. 3.
@@ III.2 QKD Performance Under Noisy Conditions
Using Eq. (18) and (14), we examine SKR and QBER in noisy scenarios. With a pair rate ($B$) of 500K cps distributed equally at the Alice and Bob sites, Fig. 6 illustrates the trend of SKR and QBER in various noise levels for different start-bin and stop-bin choices. In this demonstration, we considered equal noise distributed to all detectors. Our model also allows for unbalanced noise sharing scenarios. The result shows that a QBER threshold of $Q=0.102$ (indicated by the dark dotted line in Fig. 6) is established, above which no secure key can be generated. At this threshold, the secure key rate becomes zero ($R=0$). Therefore, achieving $Q<0.102$ is a target for the implementation of QKD. To reach this threshold, different start and stop bin choices tolerate different levels of noise. For example, when the start bin = 3 ns and the stop bin = 6 ns, the system tolerates 26 million noise counts before reaching $R=0$. However, it accepts 17 million noise counts when the start of the bin = 1 ns and the stop bin = 8 nm, despite the larger total coincidence window. When the start bin is set near the centroid of the coincidence window (the start bin = 4.5 ns) and the stop bin = 7.5 ns, the system shows reduced resistance to noise, tolerating only 13.5 million noise counts.
@@ IV Conclusion
We also presented a test to characterize detector functional changes owing to high photon rates. An experimental test to understand detector functional changes in a pulsed laser-driven QKD has previously shown [21], where changes in detector parameters such as dead time, variation in jitter, and twilight effect were investigated. Contrary to CW-QKD, pulsed lasers offer full control of the pulse time interval, including the adjustment of undesirable detection timing shifts. This would not be straightforward in a CW-driven QKD, particularly because this effect can adversely affect performance if detection rates are imbalanced. In this scenario, our model will be helpful for analysis and optimization.

### v2

@@ II.2 Coincidence Measurement in Noisy Settings
    The height of the coincidence histogram started reducing beyond detection threshold of 4.5 Mcps
@@
where $B$ is the number of pairs generated, $\eta$ is the detection efficiency. $\eta$ is inclusive of all loss parameters such as transmission loss, detector efficiency, coupling efficiency. With this, the probability of simultaneously arriving photons in both detectors can be estimated with
@@
$S_{B}=B\eta_{A}\eta_{B}+N_{B}+DC_{B},$ (4)
where $N_{A}$ and $N_{B}$ are the count of noise entered in detector $A$ and detector $B$, $DC_{A}$ and $DC_{B}$ are the dark count rates of detector $A$ and detector $B$, respectively. In the measurement, there is a random probability that two uncorrelated photon events fall within the same time window, even though they are not from the SPDC process. This is known as accidental coincidence events. These events are the primary source of noise in coincidence-based measurement. Therefore, the rate of accidental coincidence increases with the noise photons. With this, the accidental coincidence is calculated by multiplying single events from two detectors with the chosen coincidence time window which can be written as
@@
where $\sigma$ is the detector jitter and $\tau$ is the measurement time. This method can be applied in numerical simulations to investigate quantum state tomography of entangled qubits with measurements affected by time uncertainty [27]. However, this approximation fails when systems operate in extreme noise regimes. Also, this approximation does not allow one to study individual detector characteristics from the coincidence histogram. In our experiment, where only one detector receives noise-induced clicks, the probabilistic distribution of the coincidence events does not show a clean Gaussian profile in the high noise regime, rendering a normal distribution that does not fit our collected data. Neumann et al. [20] investigated coincidence events using the normal distribution function for various coincidence values, accounting for variations in the Full Width at Half Maximum (FWHM) and shifts caused by dispersion in optical fibers. Using the detector function of the form of Eq. (6), they characterized the changes in the detector parameters using linear functions. In contrast, our observations reveal a polynomial growth in FWHM (see Fig 2b) and a shift in the coincidence time. This discrepancy arises because previous work assumed equal photon pair detection rates across all detectors in a low-noise scenario (in our case, if noise photons entered <4.5 Mcps, the widening of FWHM will appear almost linear). Consequently, in their study, FWHM changes and coincidence peak shifts appeared linear.
Our results show noise photon can be responsible for shift in the detection timing, particularly its evident when detection rates are imbalanced between detectors. To address this, we should be able to cater to the changes in individual detectors reflect in the theoretical fit. By incorporating noise-induced variation parameter with exponential distribution coincidence histogram, we could detect changes in the individual detectors as follows
@@
where $W_{s(i)}$ is a coefficient used to accommodate the change in the width of coincidence events, $t_{d}$ is the dead time of the detector and $C$ is the centroid value of the coincidence histogram. $1/W$ will provide the FWHM in the time domain. By convoluting detector function in Eq. (6) with $A_{i}$ and $A_{j}$, we can obtain a complete coincidence histogram function of the form:
@@
where $\sigma_{s}$ and $\sigma_{i}$ are the timing uncertainty (jitter) of the detectors used at signal and idler arms, respectively, $t_{D}$ is the shift of centroid point of coincidence histogram from the origin, and “$\mathrm{erfc}$” represents the complimentary error function. Detector timing jitter is predominantly responsible for the width of the coincidence events. The timing uncertainty of the detectors (jitter) is responsible for the width of the coincidence histogram plots. If the temporal coherence of the collected photons is significantly greater than the jitter of the detection system, the measured spectral profile will reflect the true spectral characteristics of the photons more closely [28, 29]. The coherence time of our photon pairs is about 21.5 fs (corresponding to the spectral bandwidth of 30 nm), which is negligible and overshadowed by the timing jitter of the detectors. Therefore, timing uncertainty due to this factor could be omitted. Additionally, the temporal resolution of the timestamp unit also contributes to the width of coincidence events if it is greater than the timing jitter. Since the temporal resolution of our timestamp unit is 81 ps, we can omit this factor as well. When the noise is low, we can approximate $W_{s(i)}=1/\sigma_{s(i)}$. As noise increases, we adjusted the value of $W_{s(i)}$ to fit with the experimental data. Theoretical fit for the collected experimental data is shown in Fig.  2(a). The change in the FWHM in different noise settings is shown in Fig. 2(b). We observe that $W_{s(i)}$ is proportional to the FWHM of the coincidence histogram. The experimentally obtained shift of the start-stop histogram was directly incorporated into Eq. (9) during data fitting. One parameter that largely depends on the shape of the coincidence histogram is the variation in the dead time of the detectors due to high photon detection events. We have investigated the change in dead time with increased event rates in the detector. We used experimentally obtained $t_{d}$ values (discussed in Section II.3) in our model.
@@ II.3 Characterization of Detector Parameters
The SPCM-AQRH-10 detector is a commonly used single-photon detection module in QKD. This is an actively quenched Geiger mode detector, and the manufacturer-provided parameters of this detector are summarized in Table. 1. To study the count rate-dependent detector parameters, from the experimental setup described in Fig. 1, we only considered the APD2 output and investigated detector parameters such as detector dead time and twilight events. For this, the singles values recorded from APD2 were subtracted from their neighboring photon time values, that is, $\Delta t=t_{n+1}-t_{n}$ where $t$ is the time of arrival of photons and $n$ is the index value. The number of measured time separations ($\Delta$) is plotted as a histogram in Fig. 3. This figure gives interesting results about many detector parameters, including dead-time and twilight events. The dead time in single-photon avalanche photo detectors [21] is often described as a temporal window in which photons cannot be detected, which means that the dead time can be estimated by knowing the separation from 0 to the offset value of $\Delta t$ time in Fig. 3.
@@ III.2 QKD Performance Under Noisy Conditions
Using Eq. (18) and (14), we examine SKR and QBER in noisy scenarios. With a pair rate ($B$) of 500K cps distributed equally at the Alice and Bob sites, Fig. 6 illustrates the trend of SKR and QBER in various noise levels for different start-bin and stop-bin choices. In this demonstration, we considered equal noise distributed to all detectors. The result shows that a QBER threshold of $Q=0.102$ (indicated by the dark dotted line in Fig. 6) is established, above which no secure key can be generated. At this threshold, the secure key rate becomes zero ($R=0$). Therefore, achieving $Q<0.102$ is a target for the implementation of QKD. To reach this threshold, different start and stop bin choices tolerate different levels of noise. For example, when the start bin = 3 ns and the stop bin = 6 ns, the system tolerates 26 million noise counts before reaching $R=0$. However, it accepts 17 million noise counts when the start of the bin = 1 ns and the stop bin = 8 nm, despite the larger total coincidence window. When the start bin is set near the centroid of the coincidence window (the start bin = 4.5 ns) and the stop bin = 7.5 ns, the system shows reduced resistance to noise, tolerating only 13.5 million noise counts.
@@ IV Conclusion
We also presented a test to characterize detector functional changes owing to high photon rates. An experimental test to understand detector functional changes in a pulsed laser-driven QKD has previously been shown [21], where changes in detector parameters such as dead time, variation in jitter, and twilight effect were investigated. Contrary to CW-QKD, pulsed lasers offer full control of the pulse time interval, including the adjustment of undesirable detection timing shifts. This would not be straightforward in a CW-driven QKD, particularly because this effect can adversely affect performance if detection rates are imbalanced. In this scenario, our model will be helpful for analysis and optimization.